A metal oxide is reduced by heating it in a stream of hydrogen.It is found that after compelete reduction 3.15g of the oxide have yielded 1.05g of the metal.We may conclude that

Grade:12

3 Answers

Ramesh V
14 years ago

consider +1 and +2 oxidation states for metals

case 1 :

M2O + H2  -- > 2M  +H2O

2x+16                 2x

so, its , 2*3.15/(2x+16) = 1.05/x

 or x= 4 i.e., the compound may be Helium(but helium being inert gas doesn't for oxide)

case 2 :

MO + H2  -- > M  +H2O

x+16                 x

so, its , 3.15/(x+16) = 1.05/x

 or x = 8 (but there is no element with atomic weight of 8)

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it
here and we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE
preparation. All the best.

Regards,
Ramesh
 

sakshi
6 years ago
actually it can be much more easier . we all know that there is no metal with atomic mass of 8 and helium doesn’t form oxide . so the answer can be calculated by the following method:
mass of metal/massof oxygen(both given in ques.)= eq. mass of metal (E) /eq.mass of oxygen (i.e.8)
1.05 / (3.15-1.05) =E/8
E=4
 
Atul Patel
5 years ago
Simply apply the relationship between equivalent wt. And mass of element i.e.Mass of element/eq.of element=mass of oxid